### v1

@@ Abstract
We introduce AutoGraph, a novel autoregressive framework for generating large attributed graphs using decoder-only transformers. At the core of our approach is a reversible “flattening” process that transforms graphs into random sequences. By sampling and learning from these sequences, AutoGraph enables transformers to model and generate complex graph structures in a manner akin to natural language. In contrast to diffusion models that rely on computationally intensive node features, our approach operates exclusively on these sequences. The sampling complexity and sequence length scale linearly with the number of edges, making AutoGraph highly scalable for generating large sparse graphs. Empirically, AutoGraph achieves state-of-the-art performance across diverse synthetic and molecular graph generation benchmarks, while delivering a 100-fold generation and a 3-fold training speedup compared to leading diffusion models. Additionally, it demonstrates promising transfer capabilities and supports substructure-conditioned generation without additional fine-tuning. By extending language modeling techniques to graph generation, this work paves the way for developing graph foundation models.
@@ 1 Introduction
Recent advancements in deep generative models have revolutionized various domains of artificial intelligence, demonstrating remarkable capabilities in generating complex data types such as images (Rombach et al., 2022), natural language (Brown et al., 2020; Touvron et al., 2023a; b), and audio (Dhariwal et al., 2020; Huang et al., 2024). These achievements have been primarily driven by the development of advanced architectures or methods such as transformers and diffusion models, alongside increasingly large-scale data resources. However, the generation of graph-structured data, which is fundamental to numerous scientific applications including drug discovery (Vignac et al., 2023; Lim et al., 2020), protein design (Ingraham et al., 2019), and program synthesis (Brockschmidt et al., 2019), remains a significant challenge. This disparity primarily stems from the inherent complexity of preserving structural validity, maintaining invariance properties within graphs, and achieving scalability in real-world graph generation tasks.
To this end, diffusion-based models have emerged as a promising direction for graph generation, demonstrating effectiveness in synthesizing both classic unattributed graphs and molecules (Jo et al., 2022; Vignac et al., 2023). These approaches typically implement a denoising process in discrete graph space, simultaneously predicting edge connectivity and attributes. Yet, their practical applications are constrained by fundamental scalability limitations. The requirement for full adjacency matrix operations imposes quadratic memory complexity with respect to the number of nodes. Moreover, computing additional features in each denoising step such as spectral features, often involving cubic complexity, further increases the computational overhead.
Autoregressive approaches represent an alternative paradigm, constructing graphs sequentially by generating nodes and edges in a step-by-step manner (Liao et al., 2019; You et al., 2018). These models have demonstrated strong performance in generating small to medium-sized graphs by leveraging their ability to maintain structural validity through the generation process. Nevertheless, these models face inherent limitations: they require specialized architectures, primarily based on recurrent neural networks, to process the complex ad-hoc sequential representations of graphs, preventing them from directly leveraging the remarkable advances in large language models (LLMs). Moreover, these specialized architectures often struggle with long-range dependencies and ensuring global structural consistency, leading to significantly inferior performance compared to recent diffusion models (Vignac et al., 2023). This architectural constraint not only limits their scalability but also creates a growing performance gap as general-purpose LLMs continue to advance rapidly.
In light of these challenges, we introduce a novel paradigm that bridges the gap between graph generation and LLMs through a graph-to-sequence transformation. Our approach fundamentally reconceptualizes graphs as sequential structures while maintaining their topological properties. Instead of requiring specialized architectures or operating on graph structures, we propose a method to linearize graphs into random sequences that encode local connectivity patterns. This transformation enables direct utilization of language models for graph generation while achieving optimally linear complexity with respect to the number of edges in both computational and memory requirements. Our approach effectively addresses the limitations of both diffusion-based and autoregressive methods: it maintains structural validity while enabling efficient scaling to large graphs and leveraging the powerful capabilities of modern language models.
Our work presents several technical contributions to the field of graph generation. (1) We introduce the concept of segmented Eulerian neighborhood trails (SENTs), a specialized class of Eulerian trails that permit breaks and incorporate neighborhood information. We establish sufficient conditions under which they can be employed for effective graph generation. (2) We propose an efficient flattening algorithm that transforms graphs into sequences and vice versa by sampling these SENTs, enabling lossless sequence representation of graphs. (3) Our method, termed AutoGraph, achieves state-of-the-art (SOTA) performance across diverse synthetic and molecular graph generation benchmarks, delivering a 100-fold generation and a 3-fold training speedup compared to diffusion-based models while maintaining the ability to scale to graphs of possibly immense size. (4) Additionally, AutoGraph demonstrates strong transfer learning capabilities and supports substructure-conditioned generation without additional fine-tuning. Our work not only advances the field of graph generation but also opens new avenues for applying LLMs to graph-centric tasks, paving the way for building foundation models for graphs.
@@ 2 Methods
Figure 1: Overview of AutoGraph: (1) We use Algorithm 1 to sample a SENT $s$ from the input graph: $s=(s_{1},s_{2})$ with $s_{1}=((v_{1},\emptyset),(v_{2},\emptyset),(v_{3},\emptyset))$ and $s_{2}=((v_{5},\{v_{2}\}),(v_{4},\emptyset))$. (2) We tokenize it by reindexing the vertices based on their first occurrence order in $s$ and adding special tokens (‘/’ represents breakage between segments, ‘$\bm{<}$’ and ‘$\bm{>}$’ indicate the start and end of a neighborhood set). (3) We perform the next token prediction on the tokenized sequences using a decoder-only transformer or any language model.
@@ 2.1 Segmented Eulerian Trail
The set of all generalized trails is denoted as ${\mathcal{T}}$, noting that ${\mathcal{T}}_{G}\subseteq{\mathcal{T}}$ for any $G$. For a generalized trail $w\in{\mathcal{T}}$, we define $V_{w}\subseteq V$ and $E_{w}\subseteq E$ as the sets of vertices and edges traversed by $w$, respectively, termed the generated sets of $w$. An Eulerian trail is a trail that visits every edge in a graph exactly once. Such trails are of particular interest as they capture the complete topology of the graph. However, the existence of an Eulerian trail depends on specific conditions related to vertex degrees and connectivity (Biggs et al., 1986). To generalize this concept to arbitrary graphs, we introduce the notion of trail segments:
@@
This isomorphism partitions ${\mathcal{S}}$ into equivalence classes. Moreover, we have the following relationship between SETs and graphs, relevant for our tokenization (Sec. 2.4): {theoremEnd}theorem For any SETs $s,t\in{\mathcal{S}}$, their generated graphs are isomorphic, i.e., $G_{s}\simeq G_{t}$, if $s\simeq t$. Conversely, if two graphs $G\simeq H$, then for any SET $s\in{\mathcal{S}}_{G}$, there exists a SET $t\in{\mathcal{S}}_{H}$ s.t. $s\simeq t$. {proofEnd} By definition of the isomorphism between $s$ and $t$, there exists a bijection $\pi:V_{s}\to V_{t}$ s.t. $\pi(s)=t$. Now if $u,v\in V_{s}$ are adjacent in $G_{s}$, i.e., $(u,v)\in E_{s}$, then $(\pi(u),\pi(v))$ is an edge visited by $\pi(s)=t$, thus $(\pi(u),\pi(v))\in E_{t}$. Similarly, the reverse is also true. Consequently, $G_{s}\simeq G_{t}$.
Now assume that $G\simeq H$ with an isomorphism $\pi$ and $s\in{\mathcal{S}}_{G}$. It is easy to show that $\pi(s)$ is also a SET and its generated graph $G_{\pi(s)}=H$. By taking $t=\pi(s)$, we obtain the result. While a SET in $G$ fully characterizes its structure, we show below the prefixes of the SET do not necessarily describe the substructures of $G$, a critical property for effective autoregressive graph generation.
@@
{theoremEnd}
lemma For any graph $G$ and SET $s$ in $G$, the generated graph of any prefix of $s$ is a subgraph of $G$, but not necessarily an induced subgraph. {proofEnd} Assume that $t$ is a prefix of $s$. Then $V_{t}\subseteq V_{s}=V_{G}$ and $E_{t}\subseteq E_{s}$. However, $G_{t}$ is not necessarily an induced subgraph of $G_{s}$. We consider the following counter-example: $s=((1,2,3,4,1,3))$, $V_{s}=\{1,2,3,4\}$, and $E_{s}=\{(1,2),(2,3),(3,4),(1,4),(1,3)\}$. Let $t=((1,2,3,4,1))$. $t$ is clearly a prefix of $s$, but its generated graph is not an induced subgraph of $G_{s}$ as its generated edge set does contain $(1,3)$. This result motivates us to extend the definition of generalized trails to incorporate the full structural information of the induced subgraphs, rather than arbitrary subgraphs, to constrain the generation space better and address long-range dependency challenges. Without this extension, dependencies between neighboring nodes may span a long sequence of generation steps, making it more difficult for the model to learn such dependencies. Empirically, we show that SET fails to accurately generate graphs in Section 4.4.
@@ 2.2 Segmented Eulerian Neighborhood Trail
To make the prefixes of a SET encode richer information, we need to extend SET to contain neighborhood information in a graph. Thus, we consider the following definition:
@@
A neighborhood sequence is a sequence of tuples $w:=(w_{0},\dots,w_{k})$ where $w_{i}=(v_{i},A_{i})$ with a node $v_{i}\in V$ and a neighborhood set $A_{i}\subseteq V$, $\forall i\in\{0,\dots,k\}$. $w$ is called Hamiltonian if its node sequence $n(w):=(v_{0},\dots,v_{k})$ has non-repeated elements. Moreover, $w$ is called causal if $A_{i}$ only contains visited nodes, i.e., $A_{i}\subseteq\{v_{0},\dots,v_{i-1}\}$ $\forall i\in[k]$.
We now extend this to trails:
@@
Similarly to SETs, the generated graph of a SENT $s$ is denoted by $G_{s}=(V_{s},E_{s})$. If a graph $G\simeq G_{s}$, we say that $s$ is a SENT in $G$. We denote by $\mathcal{S}^{\mathcal{N}}$ and $\mathcal{S}^{\mathcal{N}}_{G}$ the set of SENTs and SENTs in $G$. Analogously to SETs, we define an isomorphism over $\mathcal{S}^{\mathcal{N}}$ and obtain the same relationship as in Thm. 2.4. A prefix of a SENT is defined similarly to that of a SET. We give below conditions to force generated graphs of prefixes of a SENT to be induced subgraphs.
@@
{theoremEnd}
theorem For any causal SENT $s\in\mathcal{S}^{\mathcal{N}}$, the generated graph of any prefix $t$ of $s$ is an induced subgraph of $G_{s}$ if and only if $s$ is semi-hamiltonian. In this case, $s$ is called subgraph-induced. {proofEnd} Let us first introduce some notations. We denote by $R_{s}$ the sequence of the start tuples across all neighborhood trails in $s$, which is also a neighborhood sequence. By definition of semi-hamiltonian, the occurrences after the first time of a node in $s$ should be in $R_{s}$. We denote by $n(s)$ the associated node sequence of SENT $s$, i.e., $n(s):=n(\mathbin{\|}s)$.
Let us first assume that $s$ is semi-hamiltonian.
Assume that $t$ is a prefix of $s$. It is easy to show that $G_{t}$ is a subgraph of $G_{s}$. Now assume that $u,v\in V_{t}$ s.t. $(u,v)\in E_{s}$, we want to show that $(u,v)\in E_{t}$. There are two cases:
1) Assume that $u,v\in n(t)$. Since $s$ is semi-hamiltonian, $n(s)\setminus n(t)$ either does not contain $u$ or $v$, or even if one of them, say $u\in n(s)\setminus n(t)$, we have $u\in n(R_{s})$ and its associated neighborhood set is empty. In both cases, the edge $(u,v)$ does not belong to the generated edge set of the neighborhood subsequence after $\mathbin{\|}t$. By the disjointness of the generated edge sets of $s$, it can only be included in the generated edge set of $t$, we thus have $(u,v)\in E_{t}$.
2) Assume that one of them, say $u\notin n(t)$. There exists a neighborhood set $A$ in a tuple of $\mathbin{\|}t$ such as $u\in A$. Since $t$ is causal, we have $u\in n(t)$ which contradicts the assumption.
In all the above cases, we have $(u,v)\in E_{t}$.
Now let us assume that the generated graph of any prefix of $s$ is an induced subgraph of $G_{s}$.
Let us prove that $s$ is semi-hamiltonian by contradiction. Assume that there exist two tuples in $\mathbin{\|}s$ with the same nodes $s_{i}=(v,A_{i})$ and $s_{j}=(v,A_{j})$ with $i<j$. There are two cases: 1) $s_{j}\notin R_{s}$. A tuple $(u,A_{u})$ exists one step before $s_{j}$ in the same neighborhood trail. We consider the prefix $t$ ending at $(u,A_{u})$. We have $v,u\in V_{t}$ and $(u,v)\in E_{s}$, but $(u,v)\notin E_{t}$, by the disjointness of $s$ and since $(u,v)$ is visited at $s_{j}$ after $t$. 2) $s_{j}\notin R_{s}$ and $A_{j}\neq\emptyset$. Since $s$ is causal, there exists $s_{u}:=(u,A_{u})$ before $s_{j}$ s.t. $u\in A_{j}$. We consider the prefix $t$ ending at exactly this tuple. We have $u,v\in V_{t}$ and $(u,v)\in E_{s}$, but $(u,v)\notin E_{t}$, by the disjointness of $s$ and since $(u,v)$ is an edge visited at $(v,A_{j})$ after $t$. Now let us find the conditions for a causal and Hamiltonian SENT. For any SENT $s$ and a tuple $w:=(v,A)$ in $s$, we denote by $V_{s}(w)$ the set of nodes visited by $s$ before $w$, excluding the node linked to $v$ through the trail if it exists. We have the following necessary and sufficient conditions: {theoremEnd}theorem For $s\in\mathcal{S}^{\mathcal{N}}_{G}$, $s$ is causal and Hamiltonian if and only if every tuple $w:=(v,A_{v})$ in $\mathbin{\|}s$ satisfies $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(w)$. In this case, every node is visited exactly once. Moreover, $s$ is causal and semi-hamiltonian if and only if every tuple $w:=(v,A_{v})$ in $s$ satisfies either $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(w)$ or $A_{v}=\emptyset$. {proofEnd} Let us first assume that for any tuple $w:=(v,A_{v})$ in $\mathbin{\|}s$, $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(w)$. Since $A_{v}\subseteq V_{s}(w)$ which is a subset of the set of visited nodes, $s$ is causal. Now we prove $s$ is Hamiltonian by contradiction. Assume that there exist two tuples in $\mathbin{\|}s$, $s_{u}:=(u,A_{u})$ and a later visited one $s_{v}:=(v,A_{v})$ s.t. $u=v$. Then, $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(s_{v})={\mathcal{N}}_{G}(u)\cap V_{s}(s_{v})$ should contain the node visited before that is a neighbor of $u$ (either through a trail or the neighborhood set of $u$), denoted by $u^{\prime}$. Thus, the edge $(u,u^{\prime})$ has been visited twice, which contradicts the disjointness of $s$.
Assuming that $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(w)$ or $A_{v}=\emptyset$ for any tuple $(v,A_{v})$ in $s$, we can also prove $s$ is semi-hamiltonian by contradiction. Assume that there exist two tuples in $\mathbin{\|}s$, $s_{u}:=(u,A_{u})$ and a later visited one $s_{v}:=(v,A_{v})$ s.t. $u=v$ and $A_{v}\neq\emptyset$. Then, $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(s_{v})={\mathcal{N}}_{G}(u)\cap V_{s}(s_{v})$ by assumption. And using the same argument as above, we have the contradiction.
Now assume that $s$ is causal and Hamiltonian. Let us prove the other direction by contradiction. There exists a tuple $w:=(v,A_{v})$ in $s$ s.t. $A_{v}\neq{\mathcal{N}}_{G}(v)\cap V_{s}(w)$. As $s$ is causal, $A_{v}\subseteq V_{s}(w)$. $A_{v}\subseteq{\mathcal{N}}_{G}(v)$ as $s\in\mathcal{S}^{\mathcal{N}}_{G}$. Thus, $A_{v}\subset{\mathcal{N}}_{G}(v)\cap V_{s}(w)$, which means that there exists $u\in{\mathcal{N}}_{G}(v)\cap V_{s}(w)$ and $u\notin A_{v}$. Hence, $(u,v)\in E_{G}$ and $u$ is visited before $v$. However, as $u\notin A_{v}$, $(u,v)\in E_{G}$, and $s$ is Hamiltonian, there exists a tuple $(u,A_{u})$ in $\mathbin{\|}s$ s.t. $v\in A_{u}$. By causality of $s$, $v$ is visited before $u$, which contradicts the fact that $s$ is Hamiltonian.
Assuming that $s$ is causal and semi-hamiltonian. Let us prove the other direction by contradiction. There exists a tuple $w:=(v,A_{v})$ in $s$ s.t. $A_{v}\neq{\mathcal{N}}_{G}(v)\cap V_{s}(w)$ and $A_{v}\neq\emptyset$. Using the same arguments as above, there exists $(u,v)\in E_{G}$, and $u$ is visited before $v$. However, as $u\notin A_{v}$ and $(u,v)\in E_{G}$, $s$ should visit the edge $(u,v)$ at some point. Since $s$ is semi-hamiltonian, if $s$ visits again $u,v$ they can only be the first nodes and their associated neighborhood sets are empty. Hence, there is no means for $s$ to visit $(u,v)$ after $v$, leading to contradiction. This theorem offers a simple sufficient condition for subgraph-induced SENTs. We provide in the following an implementation through a random path sampling strategy.

### v2

@@ Abstract
We introduce AutoGraph, a scalable autoregressive model for attributed graph generation using decoder-only transformers. By flattening graphs into random sequences of tokens through a reversible process, AutoGraph enables modeling graphs as sequences without relying on additional node features that are expensive to compute, in contrast to diffusion-based approaches. This results in sampling complexity and sequence lengths that scale optimally linearly with the number of edges, making it scalable and efficient for large, sparse graphs. A key success factor of AutoGraph is that its sequence prefixes represent induced subgraphs, creating a direct link to sub-sentences in language modeling. Empirically, AutoGraph achieves state-of-the-art performance on synthetic and molecular benchmarks, with up to 100x faster generation and 3x faster training than leading diffusion models. It also supports substructure-conditioned generation without fine-tuning and shows promising transferability, bridging language modeling and graph generation to lay the groundwork for graph foundation models.
Our code is available at https://github.com/BorgwardtLab/AutoGraph.
@@ 1 Introduction
Recent advancements in deep generative models have revolutionized various domains of artificial intelligence, demonstrating remarkable capabilities in generating complex data types such as images [50], natural language [7, 58, 59], and audio [15, 25]. These achievements have been primarily driven by the development of advanced architectures or methods such as transformers and diffusion models, alongside increasingly large-scale data resources. However, the generation of graph-structured data, which is fundamental to numerous scientific applications including drug discovery [61, 40], protein design [26], and program synthesis [5], remains a significant challenge. This disparity primarily stems from the inherent complexity of preserving structural validity, maintaining invariance properties within graphs, and achieving scalability in real-world graph generation tasks.To this end, diffusion-based models have emerged as a promising direction for graph generation, demonstrating effectiveness in synthesizing both classic unattributed graphs and molecules [31, 61]. These approaches typically implement a denoising process in discrete graph space, simultaneously predicting edge connectivity and attributes. Yet, their practical applications are constrained by fundamental scalability limitations. The requirement for full adjacency matrix operations imposes quadratic memory complexity with respect to the number of nodes. Moreover, computing additional node features in each denoising step, such as spectral features, often involving cubic complexity, further increases the computational overhead.Autoregressive approaches represent an alternative paradigm, constructing graphs sequentially by generating nodes and edges in a step-by-step manner [39, 66]. These models have demonstrated strong performance in generating small to medium-sized graphs by leveraging their ability to maintain structural validity through the generation process. Nevertheless, these models face inherent limitations: their sequences are not composed of tokens and thereby require specialized architectures, primarily based on recurrent neural networks, to process their complex ad-hoc sequential representations, preventing them from directly leveraging the remarkable advances in large language models (LLMs). Moreover, these specialized architectures often struggle with long-range dependencies and global structural consistency, leading to significantly inferior performance compared to recent diffusion models [61, 32]. This representational and architectural constraint not only limits their scalability but also creates a growing performance gap as general-purpose LLMs continue to advance rapidly.In light of these challenges, we introduce a novel paradigm that bridges the gap between graph generation and LLMs through a graph-to-sequence transformation. Our approach advances previous random walk-based methods by representing graphs as sequences of tokens while maintaining their topological properties. Instead of requiring specialized architectures or operating directly on graph structures, we propose a method to linearize graphs into random sequences that encode local connectivity patterns. This transformation enables direct utilization of language models for graph generation while achieving optimally linear complexity with respect to the number of edges in both computational and memory requirements. Our approach effectively addresses the limitations of both diffusion-based and autoregressive methods: it maintains structural validity while enabling efficient scaling to large graphs and leveraging the powerful capabilities of modern language models.Our work presents several technical contributions to the field of graph generation. (1) We introduce the concept of segmented Eulerian neighborhood trails (SENTs), a specialized class of Eulerian trails that permit breaks and incorporate neighborhood information. We establish sufficient conditions under which they can be employed for effective graph generation. (2) We propose an efficient flattening algorithm that transforms graphs into sequences and vice versa by sampling these SENTs, enabling lossless sequence representation of graphs. (3) Our method, termed AutoGraph, achieves state-of-the-art (SOTA) performance across diverse synthetic and molecular graph generation benchmarks, delivering a 100-fold generation and a 3-fold training speedup compared to diffusion-based models while maintaining the ability to scale to graphs of possibly immense size. (4) Additionally, AutoGraph demonstrates strong transfer learning capabilities and supports substructure-conditioned generation without additional fine-tuning. Our work not only advances the field of graph generation but also opens new avenues for applying LLMs to graph-centric tasks, paving the way for building foundation models for graphs.
@@ 2.1 Segmented Eulerian Trail
The set of all generalized trails is denoted as $\mathcal{T}$}, noting that ${\mathcal{T}}_{G}\subseteq{\mathcal{T}}$ for any $G$. For a generalized trail $w\in{\mathcal{T}}$, we define $V_{w}\subseteq V$ and $E_{w}\subseteq E$ as the sets of vertices and edges traversed by $w$, respectively, termed the generated sets of $w$.An Eulerian trail is a trail that visits every edge in a graph exactly once. Such trails are of particular interest as they capture the complete topology of the graph. However, the existence of an Eulerian trail depends on specific conditions related to vertex degrees and connectivity [3]. To generalize this concept to arbitrary graphs, we introduce the notion of trail segments:
@@
This isomorphism partitions $\mathcal{S}$} into equivalence classes. Moreover, we have the following relationship between SETs and graphs, relevant for our tokenization (Sec. 2.4): neurips_2025-pratenddefaultcategory.tex
Theorem 2.5.
For any SETs $s,t\in{\mathcal{S}}$, their generated graphs are isomorphic, i.e., $G_{s}\simeq G_{t}$, if $s\simeq t$. Conversely, if two graphs $G\simeq H$, then for any SET $s\in{\mathcal{S}}_{G}$, there exists a SET $t\in{\mathcal{S}}_{H}$ s.t. $s\simeq t$.
See proof on page LABEL:proof:prAtEndii.neurips_2025-pratenddefaultcategory.texWhile a SET in $G$ fully characterizes its structure, we show below the prefixes of the SET do not necessarily describe the substructures of $G$, a critical property for effective autoregressive graph generation.
@@
Lemma 2.8.
For any graph $G$ and SET $s$ in $G$, the generated graph of any prefix of $s$ is a subgraph of $G$, but not necessarily an induced subgraph.
See proof on page LABEL:proof:prAtEndiii.neurips_2025-pratenddefaultcategory.texThis result motivates us to extend the definition of generalized trails to incorporate the full structural information of the induced subgraphs, rather than arbitrary subgraphs, to constrain the generation space better and address long-range dependency challenges.Without this extension, dependencies between neighboring nodes may span a long sequence of generation steps, making it more difficult for the model to learn such dependencies. Empirically, we show that SET fails to generate structurally valid graphs in Section 4.4.
@@ 2.2 Segmented Eulerian Neighborhood Trail
To make the prefixes of a SET encode richer information, we need to extend SET to contain neighborhood information in a graph. Thus, we consider the following definitions:
@@
A neighborhood sequence is a sequence of tuples $w:=(w_{0},\dots,w_{k})$ where $w_{i}=(v_{i},A_{i})$ with a node $v_{i}\in V$ and a neighborhood set $A_{i}\subseteq V$, $\forall i\in\{0,\dots,k\}$. $w$ is called Hamiltonian if its node sequence $n(w):=(v_{0},\dots,v_{k})$ has non-repeated elements. $w$ is called causal if $A_{i}$ only contains visited nodes, i.e., $A_{i}\subseteq\{v_{0},\dots,v_{i-1}\}$ $\forall i\in[k]$.
@@
Similarly to SETs, the generated graph of a SENT $s$ is denoted by $G_{s}=(V_{s},E_{s})$. If a graph $G\simeq G_{s}$, we say that $s$ is a SENT in $G$. We denote by $\mathcal{S}^{\mathcal{N}}$ and $\mathcal{S}^{\mathcal{N}}_{G}$ the set of SENTs and SENTs in $G$. Analogously to SETs, we define an isomorphism over $\mathcal{S}^{\mathcal{N}}$ and obtain the same relationship as in Thm. 2.5. A prefix of a SENT is defined similarly to that of a SET. We give below conditions to force generated graphs of prefixes of a SENT to be induced subgraphs.
@@
Theorem 2.14.
For any causal SENT $s\in\mathcal{S}^{\mathcal{N}}$, the generated graph of any prefix $t$ of $s$ is an induced subgraph of $G_{s}$ if and only if $s$ is semi-hamiltonian. In this case, $s$ is called subgraph-induced.
See proof on page LABEL:proof:prAtEndiv.neurips_2025-pratenddefaultcategory.texNow let us find the conditions for a causal and Hamiltonian SENT. For any SENT $s$ and a tuple $w:=(v,A)$ in $s$, we denote by $V_{s}(w)$ the set of nodes visited by $s$ before $w$, excluding the node linked to $v$ through the trail if it exists. We have the following necessary and sufficient conditions:
Theorem 2.15.
For $\in\mathcal{S}^{\mathcal{N}}_{G}, ss$ is causal and Hamiltonian if and only if every tuple $w:=(v,A_{v})$ in $\mathbin{\|}s$ satisfies $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(w)$. In this case, every node is visited exactly once. Moreover, $s$ is causal and semi-hamiltonian if and only if every tuple $w:=(v,A_{v})$ in $s$ satisfies either $A_{v}={\mathcal{N}}_{G}(v)\cap V_{s}(w)$ or $A_{v}=\emptyset$.
See proof on page LABEL:proof:prAtEndv.neurips_2025-pratenddefaultcategory.texThis theorem offers a simple sufficient condition for subgraph-induced SENTs. We provide an implementation in the following through a random path sampling strategy.